A grocer mixes peanuts that cost $1.43 per pound and walnuts that cost $2.63 per pound to make 100 pounds of a mixture that costs $2.15 per pound. How much of each kind of nut is put into the mixture?
Answer by ankor@dixie-net.com(22740)

A grocer mixes peanuts that cost $1.43 per pound and walnuts that cost $2.63 per pound to make 100 pounds of a mixture that costs $2.15 per pound.
How much of each kind of nut is put into the mixture?
:
let x = the amt of walnuts required
The total of the final mixture is to be 100 lb, therefore
(100-x) = amt of peanuts required
:
A typical mixture equation
2.63x + 1.43(100-x) = 2.15(100)
2.63x + 143 - 1.43x = 215
2.63x - 1.43x = 215 - 143
1.20x = 72
x = 72/1.2
x = 60 lb of walnuts
then
100 - 60 = 40 lb of peanuts